# A Priori Uncertainty Quantification of Reacting Turbulence Closure Models using Bayesian Neural Networks

### v1

@@ 2.3.1 Specification of the Prior
Bayesian inference provides a way to integrate empirical data with prior information about phenomenon being modeled. In practice, the prior drives the first term in the ELBO objective function (Eq. 6), and thus the selection of the prior has a direct impact on the quality of the resultant model izmailov2021bayesian . In extreme cases, prior misspecification can only be overcome in the asymptotic limit of data, with the resultant model strictly adhering to the prior belief.
There are two ways of viewing BNN priors: the weight-space view and the functional-space view. Since BNN s operate on a parameterized distribution of the weights, $q(\mathbf{w}|\theta)$, a prior is inherently defined on the BNN parameters. However, the effect of a single weight on the output is difficult to gauge due to interactions with other weights and nonlinear layers. In other words, the lack of interpretability of the neural network parameter prevents one from meaningfully assigning a prior to the weights. In this case, using a non-informative prior such as an isotropic Gaussian could be envisioned, however, this has been shown to be sub-optimal fortuin2021bayesian . Often one does have intuition about the outputs and many methods have been derived to “tune” priors empirically. These include warm-start methods ash2020warm , initial fitting of the model to Gaussian process realizations flam2017mapping , and initialization from a trained deterministic model krishnan2020specifying . One may be tempted to instead ignore the contribution of the prior altogether, however, it has been shown that direct minimization of the negative log-likelihood is inappropriate for BNN s wei2022performance . In this work, we employ a “trainable” prior that updates the prior mean every epoch while while keeping the variance fixed. This choice changes the interpretation of the KL term to that of a data-driven regularization and has connections to the KL annealing method in bowman2015generating .
@@ 2.3.2 Enforcing Extrapolation Behavior
The aforementioned strategies provide a way to quantify epistemic and aleatoric uncertainty where data is available. Deploying data-driven models in place of physics-based models requires that the phase space spanned by the training data appropriately encompasses the phase space over which the model is queried. In general, this requirement is difficult to guarantee, which has traditionally led to adaptive methods pope1997computationally or data-driven methods complemented with synthetic data chatzopoulos2013chemistry . In the absence of training data, one would like for extrapolatory predictions to closely match those coming from physics models. Solely relying on the extrapolatory nature of neural network models is unrealistic, especially since it is still not fully understood how neural networks extrapolate xu2020neural , zhou2022domain . In the present case, if training data is not available, one may prefer that the sub-filter progress variable dissipation rate should fall back to the predictions of a physics-based model. One way to ensure this behavior would be to detect that a model is called outside of the range of the training data, in which case a separate closure model may be used. However, detecting whether or not an individual input is out-of-distribution (OOD) is a challenging problem. An alternative approach is to include data obtained from the physics-based model in the training dataset. To ensure that the synthetic data does not “pollute” the original dataset, the data is generated only in the OOD region feature space that is separate from the true data distribution, $\boldsymbol{D}$. The synthetic data can be labeled with the prediction of the physics model and an additive noise in case the physics model is subject to uncertainty.
@@
To generate OOD synthetic data, two methods are explored: a soft Brownian offset (SBO) approach moller2021out and a normalizing flow (NF) approach durkan2019neural . In the SBO approach, the OOD data is generated based on a distance metric computed with respect to all the available data. The method is particularly useful for small datasets and allows controlling the distance between the OOD region spanned and the training data. However, the distance computation can become computationally prohibitive when used with large datasets. In the NF approach, synthetic data is uniformly generated over a larger domain than the original training data. Here the data spans a hypercube whose edge size is 40% larger than the spanned amplitude of each feature. The synthetic data that overlaps with training data is subsequently discarded. To evaluate whether a synthetic data point overlaps with the original training data, the likelihood of the training dataset is estimated via a NF. Any synthetic data whose likelihood is higher than a given threshold is discarded. The method is summarized in Algorithm 1.
[…]
@@
2:Train a normalizing flow over the original data $\boldsymbol{D}$ durkan2019neural
3:Compute the probability threshold $p_{0}=min_{x_{i}\in\boldsymbol{D}}p_{NF}(x_{i})$
@@
Unlike SBO, the NF approach does not allow to control the distance between the synthetic OOD dataset and the original dataset, however, it allows fast processing of very large datasets hassanaly2023uniform . Here, the likelihood threshold $p_{0}$ is the minimal likelihood predicted over the true training set, which ensures that the OOD region spanned by the synthetic data is strictly disjoint from the original dataset $\boldsymbol{D}$.
@@
where $\mathcal{N}$ denotes a normal distribution with mean $\mu_{\rm OOD}$ being arbitrarily set to 0. More complex strategies for generating OOD labels could be formulated, such as using a linear relaxation model. An uncertainty in the OOD is set to $\sigma_{\rm OOD}=\sqrt{Var_{\boldsymbol{D}}(\chi_{\rm SFS})}$ which simplifies the implementation and the evaluation (Sec. 3.4).
@@ 3 Results and Discussion
In this section, the performance of the BNN model is demonstrated through a priori testing. As noted in Sec. 2.2, we use the dataset from Ref. yellapantula2021deep where input parameters are generated from the filtered DNS data. Model predictions using these input features are compared against filtered DNS quantities. We assess the BNN model quality on the validation dataset in Sec. 3.1. The predicted aleatoric and epistemic uncertainties and their utility for dataset augmentation are discussed in Sec. 3.2. Flame contours are modeled and studied in Sec. 3.3. Finally, the extrapolatory behavior of the model is considered in Sec. 3.4.
@@ 3.1 Model Predictive Quality
The predictive distribution of a BNN model trained only on the dataset presented in Sec. 2.2, with no additional synthetic data, is shown in Fig. 5 and Fig. 6. The mean of the model predictions on the unseen validation data are compared to the true progress variable progress variable dissipation rate as computed by the DNS in Fig. 5. An excellent agreement can be observed between the model and data, especially in regions with copious amounts of data. The log scale on the plot shows a vast majority of the data clustered along the one-to-one line between the mean prediction values and the true values. The model performance is degraded in the high progress variable dissipation rate regime. This is confirmed by Fig. 5(b) which shows that the filtering operation and the choice of the input features do not uniquely map to a $\chi_{SFS}$ value, especially when $\chi_{SFS}/\chi_{\rm lam}$ (where $\chi_{\rm lam}$ is the maximum of the progress variable dissipation rate in the freely propagating laminar flame) exceeds 2. The data-poor regime is associated with higher epistemic uncertainties, especially in the regions that are the furthest away from the parity line between the model and the data, particularly at $\chi_{SFS}/\chi_{\rm lam}>2$. Since $\chi_{SFS}$ values correspond to larger filter widths there are fewer samples available for training. While both the aleatoric and epistemic uncertainties grow with $\chi_{SFS}/\chi_{\rm lam}$, the epistemic uncertainty exhibits locally high values, which suggests that there locally data-poor regions are present in the input feature space. Those regions are the ones that should be augmented with additional data.
@@
Figure 5: Histogram plot in log-scale showing the mean prediction, $\mathbb{E}_{q(\mathbf{w}|\theta)}\left[\chi_{\rm SFS}\right]$, from BNN model and $\chi_{SFS}$ from DNS data. The dashed black line represents a perfect model.
@@
Figure 6: (a) Scatter plot showing mean predictions for individual data points shaded by the BNN model prediction of epistemic uncertainty of $\chi_{SFS}/\chi_{\rm lam}$ and (b) shaded by the BNN model prediction of aleatoric uncertainty of $\chi_{SFS}/\chi_{\rm lam}$.
@@
For further evaluation of the model performance, the conditional structure of the progress variable dissipation rate with respect to the input features is inspected. In particular, the performance with respect to three important physical parameters is investigated: the filtered progress variable $\widetilde{C}$, the sub-filter variance of the progress variable $\widetilde{C^{\prime\prime 2}}$, and the resolved progress variable diffusivity $\widetilde{D}_{C}$. Conditional means were computed with $250$ bins discretizing the input feature domain. Credible intervals are constructed by computing the BNN model output across the dataset for 250 realizations of $\mathbf{w}\sim q(\mathbf{w}|\theta)$ and retaining the $(1-\alpha)\times 100\%$ interquantile range of predicted trajectories. In this case, $\alpha=0.1$ corresponding to the $90^{\text{th}}$ percentile. For further discussion of the nuances of predictive, credible, and confidence intervals the reader is referred to smith2013uncertainty .
@@
The conditional mean profiles displayed in Fig. 7 demonstrate excellent agreement between the data and the model predictive mean. Moreover, any discrepancies, such as those in Fig. 6(c) are clearly contained by the predictive envelope. The total predictive uncertainty is dominated by the aleatoric uncertainty as suggested by Fig. 6. This is supported by direct comparison with the modeled epistemic uncertainty in Fig. 8. The interpretation is that most of the errors may be attributed to the filtering process and the feature selection rather than the lack of data. Albeit small, the epistemic uncertainty could also rapidly increase for a model deployed a posteriori in a chaotic system such as a turbulent combustion simulation hassanaly2019ensemble , hassanaly2019lyapunov .
The distribution of aleatoric uncertainty with respect to the progress variable (Fig. 6(b)) suggests that most of the aleatoric uncertainty is located in the burning region of the flame, while reduced uncertainty can be observed either in the fully burnt or unburnt regions. The distribution of aleatoric uncertainty is also skewed towards the high end of the progress variable distribution which may suggest that small-scale variability is mostly controlled by the end of the burning process, as also noted in Ref. hassanaly2019ensemble . The subfilter variance of the progress variable may be understood as a marker for the filter width of the LES. As the filter width increases, the aleatoric uncertainty rapidly increases which suggests the presence of small-scale variability that cannot be captured with the filtered features (Fig. 6(d)). The aleatoric uncertainty also appears to plateau at high progress variable variance $\widetilde{C^{\prime\prime 2}}$. Since $\widetilde{C^{\prime\prime 2}}$ characterizes small scale variability, it is increasingly difficult for the filtered input features to capture $\chi_{\rm SFS}$. This would result in monotonically increasing aleatoric uncertainty with respect to $\widetilde{C^{\prime\prime 2}}$. However, the plateauing behavior instead suggests that the subgrid scale variability of $\chi_{\rm SFS}$ that can be captured by the filtered input features saturates at medium values of $\widetilde{C^{\prime\prime 2}}$. Finally, the distribution of aleatoric uncertainty with respect to the filtered diffusion is a marker of the heterogeneity across the DNS cases included in the training dataset (Fig. 6(f)). The DNS with high diffusivity ($C_{1}$, $C_{Le}$, $D_{1}$ and $D_{Le}$ in Ref. yellapantula2021deep ) also exhibit the highest variability for the progress variable dissipation rate. These cases also had a higher Karlovitz number and turbulent kinetic energy.
@@ 3.2 Interpretation of Aleatoric and Epistemic Uncertainties
Unlike the aleatoric uncertainty, the epistemic uncertainty predicted by the BNN can be used to refine the dataset used to train the progress variable dissipation rate closure model. The task of data selection to improve model performance, especially in the context of uncertainty reduction, is known as optimal experimental design (OED) fedorov2010optimal or active learning aggarwal2014active and has connections to Bayesian optimization di2023active . To estimate and visualize the epistemic uncertainty, 250 realizations of the model $\mathbf{w}\sim q(\mathbf{w}|\theta)$ are generated and the sample variance estimate of the epistemic uncertainty $\operatorname{Var}(\mathbb{E}_{q(\mathbf{w}|\theta)}[\mathbf{y}|\mathbf{x},\boldsymbol{D}])$, is computed. The conditional mean of this quantity is computed with respect to the input features and is presented in Fig. 8.
The epistemic uncertainty follows the same trends as the ones observed for the aleatoric uncertainty for the conditioning done with respect to the progress variable (Fig. 7(a)) and the filtered diffusivity (Fig. 7(c)). The consistent distribution of uncertainty between the epistemic and the aleatoric is a consequence of the near-uniform phase-space sampling done as a pre-processing step (Sec. 2.2). Since the data is nearly uniformly distributed, information is spread across a larger phase space as more filtering is applied. This indicates that to mitigate epistemic uncertainty, data collection should instead over-represent high-aleatoric uncertainty regimes. The epistemic uncertainty conditioned with respect to the progress variable sub-filter variance (Fig. 7(b)) paints a different picture. While the epistemic uncertainty initially increases with higher filterwidth as does the aleatoric uncertainty, it does not plateau at the highest progress variable variance. Instead, it appears as though the highest filter width variances were undersampled in the final dataset. This may be a consequence of the inadequacy of performing uniform-in-phase-space sampling with a clustering approach as discussed in Ref. hassanaly2023uniform .

### v2

@@ 2.3.1 Specification of the Prior
Bayesian inference provides a way to integrate empirical data with prior information about phenomenon being modeled. In practice, the prior drives the first term in the ELBO objective function (Eq. 6), and thus the selection of the prior has a direct impact on the quality of the resultant model [71]. In extreme cases, prior misspecification can only be overcome in the asymptotic limit of data, with the resultant model strictly adhering to the prior belief.
There are two ways of viewing BNN priors: the weight-space view and the functional-space view. Since BNN s operate on a parameterized distribution of the weights, $q(\mathbf{w}|\theta)$, a prior is inherently defined on the BNN parameters. However, the effect of a single weight on the output is difficult to gauge due to interactions with other weights and nonlinear layers. In other words, the lack of interpretability of the neural network parameter prevents one from meaningfully assigning a prior to the weights. In this case, using a non-informative prior such as an isotropic Gaussian could be envisioned, however, this has been shown to be sub-optimal [72]. Often one does have intuition about the outputs and many methods have been derived to “tune” priors empirically. These include warm-start methods [73], initial fitting of the model to Gaussian process realizations [74], and initialization from a trained deterministic model [75]. One may be tempted to instead ignore the contribution of the prior altogether, however, it has been shown that direct minimization of the negative log-likelihood is inappropriate for BNN s [76]. In this work, we employ a “trainable” prior that updates the prior mean every epoch while while keeping the variance fixed. This choice changes the interpretation of the KL term to that of a data-driven regularization and has connections to the KL annealing method in [77].
@@ 2.3.2 Enforcing Extrapolation Behavior
The aforementioned strategies provide a way to quantify epistemic and aleatoric uncertainty where data is available. Deploying data-driven models in place of physics-based models requires that the phase space spanned by the training data appropriately encompasses the phase space over which the model is queried. In general, this requirement is difficult to guarantee, which has traditionally led to adaptive methods [78] or data-driven methods complemented with synthetic data [79]. In the absence of training data, one would like for extrapolatory predictions to closely match those coming from physics models. Solely relying on the extrapolatory nature of neural network models is unrealistic, especially since it is still not fully understood how neural networks extrapolate [80, 81]. In the present case, if training data is not available, one may prefer that the sub-filter progress variable dissipation rate should fall back to the predictions of a physics-based model. One way to ensure this behavior would be to detect that a model is called outside of the range of the training data, in which case a separate closure model may be used. However, detecting whether or not an individual input is out-of-distribution (OOD) is a challenging problem. An alternative approach is to include data obtained from the physics-based model in the training dataset. To ensure that the synthetic data does not “pollute” the original dataset, the data is generated only in the OOD region feature space that is separate from the true data distribution, $\boldsymbol{D}$. The synthetic data can be labeled with the prediction of the physics model and an additive noise in case the physics model is subject to uncertainty.
@@
To generate OOD synthetic data, two methods are explored: a soft Brownian offset (SBO) approach [82] and a normalizing flow (NF) approach [83]. In the SBO approach, the OOD data is generated based on a distance metric computed with respect to all the available data. The method is particularly useful for small datasets and allows controlling the distance between the OOD region spanned and the training data. However, the distance computation can become computationally prohibitive when used with large datasets. In the NF approach, synthetic data is uniformly generated over a larger domain than the original training data. Here the data spans a hypercube whose edge size is 40% larger than the spanned amplitude of each feature. The synthetic data that overlaps with training data is subsequently discarded. To evaluate whether a synthetic data point overlaps with the original training data, the likelihood of the training dataset is estimated via a NF. Any synthetic data whose likelihood is higher than a given threshold is discarded. The method is summarized in Algorithm 1.
Algorithm 1 Out-of-distribution data generation with normalizing flows. $p_{\rm NF}$ denotes the probability predicted by the normalizing flow and $x_{i}$ denotes the ith element of the dataset $\boldsymbol{D}$.
@@
2:Train a normalizing flow over the original data $\boldsymbol{D}$ [83]
3:Compute the probability threshold $p_{0}=\operatorname{min}_{x_{i}\in\boldsymbol{D}}p_{\rm NF}(x_{i})$
@@
Unlike SBO, the NF approach does not allow to control the distance between the synthetic OOD dataset and the original dataset, however, it allows fast processing of very large datasets [25]. Here, the likelihood threshold $p_{0}$ is the minimal likelihood predicted over the true training set, which ensures that the OOD region spanned by the synthetic data is strictly disjoint from the original dataset $\boldsymbol{D}$.
@@
where $\mathcal{N}$ denotes a normal distribution with mean $\mu_{\rm OOD}$ being arbitrarily set to 0. The uncertainty in the OOD is set to $\sigma_{\rm OOD}=\operatorname{Var}_{\boldsymbol{D}}(\chi_{\rm SFS})^{1/2}$ which simplifies the implementation and the evaluation (Sec. 3.4). More complex strategies for generating OOD labels could be formulated, such as labeling the synthetic OOD data using a linear relaxation model. Here, an arbitrary marker is used which can primarily serve to detect OOD queries (see Sec. 3.5).
@@ 3 Results and Discussion
In this section, the performance of the BNN model is demonstrated through a priori testing. As noted in Sec. 2.2, we use the dataset from Ref. [46] where input parameters are generated from the filtered DNS data. Model predictions using these input features are compared against filtered DNS quantities. We assess the BNN model quality on the validation dataset in Sec. 3.1. The predicted aleatoric and epistemic uncertainties and their utility for dataset augmentation are discussed in Sec. 3.2. Flame contours are modeled and studied in Sec. 3.3. Finally, the extrapolatory behavior of the model is considered in Sec. 3.4 and Sec. 3.5.
@@ 3.1 Model Predictive Quality
The predictive distribution of a BNN model trained only on the dataset presented in Sec. 2.2, with no additional synthetic data, is shown in Fig. 5 and Fig. 6. The mean of the model predictions on the unseen validation data are compared to the true progress variable progress variable dissipation rate as computed by the DNS in Fig. 5. An excellent agreement can be observed between the model and data, especially in regions with copious amounts of data. The log scale on the plot shows a vast majority of the data clustered along the one-to-one line between the mean prediction values and the true values. The model performance is degraded in the high progress variable dissipation rate regime. This is confirmed by Fig. 6(b) which shows that the filtering operation and the choice of the input features do not uniquely map to a $\chi_{\rm SFS}$ value, especially when $\chi_{\rm SFS}/\chi_{\rm lam}$ (where $\chi_{\rm lam}$ is the maximum of the progress variable dissipation rate in the freely propagating laminar flame) exceeds 2. The data-poor regime is associated with higher epistemic uncertainties, especially in the regions that are the furthest away from the parity line between the model and the data, particularly at $\chi_{\rm SFS}/\chi_{\rm lam}>2$. Since $\chi_{\rm SFS}$ values correspond to larger filter widths there are fewer samples available for training. While both the aleatoric and epistemic uncertainties grow with $\chi_{\rm SFS}/\chi_{\rm lam}$, the epistemic uncertainty exhibits locally high values, which suggests that there locally data-poor regions are present in the input feature space. Those regions are the ones that should be augmented with additional data.
@@
Figure 5: Histogram plot in log-scale showing the mean prediction, $\mathbb{E}_{q(\mathbf{w}|\theta)}\left[\chi_{\rm SFS}\right]$, from BNN model and $\chi_{\rm SFS}$ from DNS data. The dashed black line represents a perfect model.
@@
Figure 6: (a) Scatter plot showing mean predictions for individual data points shaded by the BNN model prediction of epistemic uncertainty of $\chi_{\rm SFS}/\chi_{\rm lam}$ and (b) shaded by the BNN model prediction of aleatoric uncertainty of $\chi_{\rm SFS}/\chi_{\rm lam}$.
@@
For further evaluation of the model performance, the conditional structure of the progress variable dissipation rate with respect to the input features is inspected. In particular, the performance with respect to three important physical parameters is investigated: the filtered progress variable $\widetilde{C}$, the sub-filter variance of the progress variable $\widetilde{C^{\prime\prime 2}}$, and the resolved progress variable diffusivity $\widetilde{D}_{C}$. Conditional means were computed with $250$ bins discretizing the input feature domain. Credible intervals are constructed by computing the BNN model output across the dataset for 250 realizations of $\mathbf{w}\sim q(\mathbf{w}|\theta)$ and retaining the $(1-\alpha)\times 100\%$ interquantile range of predicted trajectories. In this case, $\alpha=0.1$ corresponding to the $90^{\text{th}}$ percentile. For further discussion of the nuances of predictive, credible, and confidence intervals the reader is referred to [51].
@@
The conditional mean profiles displayed in Fig. 7 demonstrate excellent agreement between the data and the model predictive mean. Moreover, any discrepancies, such as those in Fig. 7(c) are clearly contained by the predictive envelope. The total predictive uncertainty is dominated by the aleatoric uncertainty as suggested by Fig. 6. This is supported by direct comparison with the modeled epistemic uncertainty in Fig. 8. The interpretation is that most of the errors may be attributed to the filtering process and the feature selection rather than the lack of data. Albeit small, the epistemic uncertainty could also rapidly increase for a model deployed a posteriori in a chaotic system such as a turbulent combustion simulation [88, 89].
The distribution of aleatoric uncertainty with respect to the progress variable (Fig. 7(b)) suggests that most of the aleatoric uncertainty is located in the burning region of the flame, while reduced uncertainty can be observed either in the fully burnt or unburnt regions. The distribution of aleatoric uncertainty is also skewed towards the high end of the progress variable distribution which may suggest that small-scale variability is mostly controlled by the end of the burning process, as also noted in Ref. [88]. The subfilter variance of the progress variable may be understood as a marker for the filter width of the LES. As the filter width increases, the aleatoric uncertainty rapidly increases which suggests the presence of small-scale variability that cannot be captured with the filtered features (Fig. 7(d)). The aleatoric uncertainty also appears to plateau at high progress variable variance $\widetilde{C^{\prime\prime 2}}$. Since $\widetilde{C^{\prime\prime 2}}$ characterizes small scale variability, it is increasingly difficult for the filtered input features to capture $\chi_{\rm SFS}$. This would result in monotonically increasing aleatoric uncertainty with respect to $\widetilde{C^{\prime\prime 2}}$. However, the plateauing behavior instead suggests that the subgrid scale variability of $\chi_{\rm SFS}$ that can be captured by the filtered input features saturates at medium values of $\widetilde{C^{\prime\prime 2}}$. Finally, the distribution of aleatoric uncertainty with respect to the filtered diffusion is a marker of the heterogeneity across the DNS cases included in the training dataset (Fig. 7(f)). The DNS with high diffusivity ($C_{1}$, $C_{Le}$, $D_{1}$ and $D_{Le}$ in Ref. [46]) also exhibit the highest variability for the progress variable dissipation rate. These cases also have a higher Karlovitz number and turbulent kinetic energy.
@@ 3.2 Interpretation of Aleatoric and Epistemic Uncertainties
Unlike the aleatoric uncertainty, the epistemic uncertainty predicted by the BNN can be used to refine the dataset used to train the progress variable dissipation rate closure model. The task of data selection to improve model performance, especially in the context of uncertainty reduction, is known as optimal experimental design (OED) [90] or active learning [91] and has connections to Bayesian optimization [92]. To estimate and visualize the epistemic uncertainty, 250 realizations of the model $\mathbf{w}\sim q(\mathbf{w}|\theta)$ are generated and the sample variance estimate of the epistemic uncertainty $\operatorname{Var}_{q(\mathbf{w}|\theta)}(\mathbb{E}[\mathbf{y}|\mathbf{x}])$, is computed. The conditional mean of this quantity is computed with respect to the input features and is presented in Fig. 8.
The epistemic uncertainty follows the same trends as the ones observed for the aleatoric uncertainty for the conditioning done with respect to the progress variable (Fig. 8(a)) and the filtered diffusivity (Fig. 8(c)). The consistent distribution of uncertainty between the epistemic and the aleatoric is a consequence of the near-uniform phase-space sampling done as a pre-processing step (Sec. 2.2). Since the data is nearly uniformly distributed, information is spread across a larger phase space as more filtering is applied. This indicates that to mitigate epistemic uncertainty, data collection should instead over-represent high-aleatoric uncertainty regimes. The epistemic uncertainty conditioned with respect to the progress variable sub-filter variance (Fig. 8(b)) paints a different picture. While the epistemic uncertainty initially increases with higher filterwidth as does the aleatoric uncertainty, it does not plateau at the highest progress variable variance. Instead, it appears as though the highest filter width variances were undersampled in the final dataset. This may be a consequence of the inadequacy of performing uniform-in-phase-space sampling with a clustering approach as discussed in Ref. [25].